OpenStudy (anonymous):

Which expression is equivalent to (1-sin^2 x)(sec x)? A) sec^2 x B)tan x C) tan^2 x D) cos x E) cos^2 x

OpenStudy (anonymous):

first try to simplify the equation to get one of the choices starting with the rule \[\sin ^{2}a+\cos ^{2}a=1\]

OpenStudy (anonymous):

How do you simplify the equation?

OpenStudy (anonymous):

1-sin^2 x=cos^2 x

OpenStudy (anonymous):

secx=1/cosx

OpenStudy (anonymous):

Is the answer cos^2 x?